# Two-generation of traceless matrices over finite fields

### v1

@@ 4.3. Sharply traceless elements
5. The exceptional case in odd characteristic
There is a single exceptional case in odd characteristic that our method with sharply traceless elements does not cover, namely the family $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F}_{q})$ with $\mathbf{F}_{q}$ of characteristic $3$. We will prove that these algebras are in fact not $2$-generated over any (finite or infinite) field of characteristic $3$.
@@ 5. The exceptional case in odd characteristic
This single failure of $2$-generation comes as a surprise, since this is a property enjoyed throughout simple algebraic structures over finite and infinite fields as pointed out in Section 1. We will show that this is a phenomenon that heavily relies both on the modularity of the situation and on smallness of the rank. However, our result diverges from [Chi18, Theorem 1], whose argument appears to hinge on the existence of a consistent matrix, but these do not exist in $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ when $\mathbf{F}$ is of characteristic $3$, see Lemma 3.1.
On the other hand, the claim that $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ is $3$-generated follows immediately from the fact that over the prime field $\mathbf{F}_{3}$ of $\mathbf{F}$, the matrices
@@
This identity implies that for any $x,y\in\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$, the vector space generated by $x,y,[x,y]$ is closed with respect to Lie bracketing with $x$ and $y$. In other words, the Lie algebra generated by $x$ and $y$ is of dimension at most $3$. Thus $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$ is not $2$-generated. Furthermore, it follows from the same identity that any two elements in $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ generate a Lie algebra of dimension at most $4$. Note that this bound is sharp, for example the Lie algebra generated by the matrices $A$ and $B$ from (4) is of dimension $4$.
@@
Appendix A Even characteristic
We have assumed throughout the paper that the characteristic of the underlying field is odd. The reason for this is that our method of proving $2$-generation relies on constructing consistent matrices, which do not exist in even characteristic.
We have performed a random computer search for generating pairs in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$ for all $n\leq 20$. Generating pairs were found in all cases apart from $n=4$ (see below). Based on this, we believe this should be the case in general.
Conjecture.
If $n\neq 4$, then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$ is $2$-generated.
The case $n=4$ is exceptional and similar to the phenomenon observed in characteristic $3$ in Section 5. We are able to prove this with the same method.
Theorem A.1.
@@ Appendix A Even characteristic
(6) $A=\begin{pmatrix}1&\cdot&\cdot&\cdot\\ \cdot&1&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\end{pmatrix},\ B=\begin{pmatrix}\cdot&\cdot&1&1\\ 1&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ 1&1&\cdot&\cdot\end{pmatrix},\ B^{\operatorname{T}}=\begin{pmatrix}\cdot&1&\cdot&1\\ \cdot&\cdot&\cdot&1\\ 1&\cdot&\cdot&\cdot\\ 1&\cdot&\cdot&\cdot\end{pmatrix}$
@@
The difficult part of the proof of Theorem A.1 again follows from an incidental identity that relies on just the right smallness of the rank and even characteristic.
Proposition A.2.
@@
For any $x,y\in\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$, the Lie algebra generated by $x$ and $y$ contains the following elements:
@@
Since Lie algebra $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ is $p$-restricted, we have $\operatorname{ad}_{z^{2}}=(\operatorname{ad}_{z})^{2}$ for any $z$. This implies that $[x,y,y,y]=[x,y,y^{2}]$ and $[x,y,x,y]=[y,x,x,y]=[y,x^{2},y]=[x^{2},y,y]=[x^{2},y^{2}]$. It then follows from Proposition A.2 that the elements (7) all belong to the vector space generated by

### v2

@@ 4.3. Sharply traceless elements
An alternative (but longer) route [^4] We thank the referee for pointing out the simpler argument using group algebras. to proving Theorem 4.5 is by mimicking the usual approach for finding a normal element, compare with [vzGG90] or [LN83] and the closely related notion of a cyclic element from [NP95].
5. Even characteristic
Our method for proving $2$-generation in the previous sections used consistent matrices. These do not exist when the field is of even characteristic. In this situation, we instead rely on semiconsistent matrices. These are diagonal matrices
$D=\operatorname{diag}(\lambda_{1},\dots,\lambda_{n})\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$
that satisfy the condition
$\lambda_{i}+\lambda_{j}=\lambda_{k}+\lambda_{l}\ \textup{ if and only if }\ \{i,j\}=\{k,l\}\text{ or }(i,k)=(j,l).$
For instance, if $\alpha\in\mathbf{F}_{q^{n}}$ is a sharply traceless element, $\operatorname{diag}(\alpha,\operatorname{Fr}(\alpha),\dots,\operatorname{Fr}^{n-1}(\alpha))$ is semiconsistent in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q^{n}})$ as long as $n\neq 4$.
Proposition 5.1.
Let $p=2$ and $n\neq 2,4$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
Proof.
It suffices to prove that $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$ is $2$-generated. Let $\alpha\in\mathbf{F}_{2^{n}}$ be a sharply traceless element with minimal polynomial $f\in\mathbf{F}_{2}[x]$ of degree $n$. Let $C\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$ be the companion matrix of $f$ with diagonal form
$D=PCP^{-1}=\operatorname{diag}\left(\alpha,\operatorname{Fr}(\alpha),\dots,\operatorname{Fr}^{n-1}(\alpha)\right)\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2^{n}}).$
Since $\alpha$ is sharply traceless, $D$ is semiconsistent.
Now let $X=(x_{ij})_{i,j}$ be an arbitrary matrix in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2^{n}})$. We have
$[X,D,\stackrel{{\scriptstyle k}}{{\dots}},D]=\sum_{i<j}(\lambda_{i}+\lambda_{j})^{k}\left(x_{ij}E_{ij}+x_{ji}E_{ji}\right),$
so arguing as in the proof of Proposition 2.1, the matrix $A_{ij}=x_{ij}E_{ij}+x_{ji}E_{ji}$ for $i\neq j$ is contained in the Lie algebra generated by $X$ and $D$.
Let us consider $X=P(E_{21}+E_{13})P^{T}$. Recall from Lemma 4.2 that the rows of $P$ are $v_{0},v_{1},\dots,v_{n-1}$, where $v_{0}=(1,\alpha,\alpha^{2},\dots,\alpha^{n-1})$ and $v_{i}=\operatorname{Fr}^{i}(v_{0})$. Thus
$e_{k}^{\operatorname{T}}PE_{ij}P^{T}e_{l}=(e_{k}^{\operatorname{T}}Pe_{i})(e_{j}^{\operatorname{T}}P^{T}e_{l})=\operatorname{Fr}^{k-1}(\alpha^{i-1})\operatorname{Fr}^{l-1}(\alpha^{j-1})$
and so
$x_{kl}=\operatorname{Fr}^{k-1}(\alpha)\operatorname{Fr}^{l-1}(1)+\operatorname{Fr}^{k-1}(1)\operatorname{Fr}^{l-1}(\alpha^{2})=\operatorname{Fr}^{k-1}(\alpha)+\operatorname{Fr}^{l}(\alpha).$
Since we are in even characteristic and $\alpha$ is sharply traceless, $x_{kl}=0$ if and only if $l\equiv k-1\pmod{n}$. In particular, $E_{n1}=x_{n1}^{-1}A_{n1}$ and $E_{i,i+1}=x_{i,i+1}^{-1}A_{i,i+1}$ for $11\leq i\leq n-$ 1 are in the Lie algebra generated by $D$ and $X.$ Since the matrices $E_{n1}$ and $E_{i,i+1}$ for $11\leq i\leq n-$ 1 generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2^{n}})$, the matrices $D$ and $X$ generate it as well, and so the matrices $C=P^{-1}DP$ and $P^{-1}XP=(E_{21}+E_{13})P^{\operatorname{T}}P$ also generate it.
Note, however, that both $C$ and $(E_{21}+E_{13})P^{\operatorname{T}}P$ belong to $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$, since
$e_{k}^{\operatorname{T}}P^{\operatorname{T}}Pe_{l}=(Pe_{k})^{\operatorname{T}}(Pe_{l})=\sum_{i=0}^{n-1}\operatorname{Fr}^{i}(\alpha^{k-1})\operatorname{Fr}^{i}(\alpha^{l-1})=\operatorname{Tr}(\alpha^{k+l-2})$
for any $k,l$. Therefore, as in the proof of Corollary 4.3, the matrices $C$ and $(E_{21}+E_{13})P^{\operatorname{T}}P$ generate the Lie algebra $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{2})$. ∎
We still have to take care of the Lie algebras of rank $n=2$ and $n=4$. Clearly $\operatorname{\mathfrak{sl}}_{2}(\mathbf{F}_{2})$ is $2$-generated, for instance by $E_{12}$ and $E_{21}$. We deal with $\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$ in Section 6.
6. Exceptional cases
There are two exceptions that our methods with sharply traceless elements do not cover, namely the families $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F}_{q})$ with $\mathbf{F}_{q}$ of characteristic $3$ and $\mathfrak{sl}_{4}(\mathbf{F}_{q})$ with $\mathbf{F}_{q}$ of characteristic $2$ We will prove that these algebras are in fact not $2$-generated, even over infinite fields of corresponding characteristic. This failure of $2$-generation comes as a surprise, since this is a property enjoyed throughout simple algebraic structures over finite and infinite fields as pointed out in Section 1. We will show that this is a phenomenon that heavily relies both on the modularity of the situation and on smallness of the rank.
Let us first deal with the case $(n,p)=(3,3)$.
@@ 6. Exceptional cases
Note that our result diverges from [Chi18, Theorem 1], whose argument appears to hinge on the existence of a consistent matrix, but these do not exist in $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ when $\mathbf{F}$ is of characteristic $3$, see Lemma 3.1.
The claim that $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ is $3$-generated follows immediately from the fact that over the prime field $\mathbf{F}_{3}$ of $\mathbf{F}$, the matrices
@@
This identity implies that for any $x,y\in\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$, the vector space spanned by $x,y,[x,y]$ is closed with respect to Lie bracketing with $x$ and $y$. In other words, the Lie algebra generated by $x$ and $y$ is of dimension at most $3$. Thus $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$ is not $2$-generated. Furthermore, it follows from the same identity that any two elements in $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ generate a Lie algebra of dimension at most $4$. Note that this bound is sharp, for example the Lie algebra generated by the matrices $A$ and $B$ from (3) is of dimension $4$.
@@
The case $(n,p)=(4,2)$ is proved with the same method.
Theorem 6.3.
@@
(5) $A=\begin{pmatrix}1&\cdot&\cdot&\cdot\\ \cdot&1&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\end{pmatrix},\ B=\begin{pmatrix}\cdot&\cdot&1&1\\ 1&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot\\ 1&1&\cdot&\cdot\end{pmatrix},\ B^{\operatorname{T}}=\begin{pmatrix}\cdot&1&\cdot&1\\ \cdot&\cdot&\cdot&1\\ 1&\cdot&\cdot&\cdot\\ 1&\cdot&\cdot&\cdot\end{pmatrix}.$
@@
The difficult part of the proof of Theorem 6.3 again follows from an incidental identity that relies on just the right smallness of the rank and even characteristic.
Proposition 6.4.
@@
For any $x,y\in\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$, the Lie algebra generated by $x$ and $y$ contains the following elements:
@@
Since the Lie algebra $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ is $2$-restricted, we have $\operatorname{ad}_{z^{2}}=(\operatorname{ad}_{z})^{2}$ for any $z$. This implies that $[x,y,y,y]=[x,y,y^{2}]$ and $[x,y,x,y]=[y,x,x,y]=[y,x^{2},y]=[x^{2},y,y]=[x^{2},y^{2}]$. It then follows from Proposition 6.4 that the elements (6) all belong to the vector space spanned by